OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 4y = 5?

OpenStudy (anonymous):

@jdoe0001

jimthompson5910 (jim_thompson5910):

hint: \[\Large b^x = y\] converts to \[\Large \log_{b}(y) = x\]

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}^y={\color{blue}{ b}}\implies log_{\color{red}{ a}}{\color{blue}{ b}}=y \\ \quad \\ \quad \\ {\color{red}{ 4}}^y = {\color{blue}{ 5}}\implies ?\)

OpenStudy (anonymous):

log_(4)5

OpenStudy (jdoe0001):

yeap =y

OpenStudy (anonymous):

what would the graph look like?

OpenStudy (jdoe0001):

well.. is a fixed value.... so... just get the constant value for "y" that is, get \(\bf log_45\)

OpenStudy (jdoe0001):

so y = "some constant" means is a horizontal line at that point

OpenStudy (anonymous):

I need the graph which is unfortunate bc I keep looking it up and its none of the options

jimthompson5910 (jim_thompson5910):

maybe you're looking for a graph that has the root of \[\Large \log_{4}(5)\]

jimthompson5910 (jim_thompson5910):

\[\Large \log_{4}(5) = \frac{\log(4)}{\log(5)} \approx 0.86135\]

OpenStudy (anonymous):

I need the graph not the answer

OpenStudy (jdoe0001):

hmmm is a constant... thus if you notice jim_thompson5910 's line above y = 0.86135 so that's the line, that's the graph

OpenStudy (jdoe0001):

hmm actaully.... would be \( \log_{4}(5) = \cfrac{\log(5)} {\log(4)}\)

jimthompson5910 (jim_thompson5910):

oh right my bad lol

jimthompson5910 (jim_thompson5910):

swapped the numbers

OpenStudy (jdoe0001):

using the log change of base rule.... then just graph it, since it's just a constant... is just a horizontal line

OpenStudy (anonymous):

that's not an option :/

OpenStudy (anonymous):

@jim_thompson5910 please help

jimthompson5910 (jim_thompson5910):

well \[\Large \log_{4}(5) = \frac{\log(5)}{\log(4)} \approx 1.160964047\]

jimthompson5910 (jim_thompson5910):

unfortunately there are 2 graphs that have a root near 1.160964047 (so it looks like)

OpenStudy (anonymous):

which one would you choose @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the two choices I mentioned both seem equally likely I have no idea which one I would choose (other than just randomly picking one)

OpenStudy (anonymous):

Did you finally got the answer because I'm stuck on it